OpenStudy (abbycross167):

Can someone please help me fix my this question? let a = 2, b = -5, c = 7, and d = -4. Evaluate b – a(cd + b)

OpenStudy (abbycross167):

My answer~ (cd+b)=-33 (7)(-4)+(-5) (7)(-4)=-28 (-28)+(-5)=-33 (-5)-(2)((-33)) (2)((-33))=-6 (-5)-(-6)=1

OpenStudy (abbycross167):

What my teacher said~ incorrect (look near the end of your work)

OpenStudy (abbycross167):

@Ineedhelplz could you try to help me?

OpenStudy (ineedhelplz):

Ok..

OpenStudy (abbycross167):

So you'll help me?

OpenStudy (ineedhelplz):

Yeah! one second

OpenStudy (ineedhelplz):

2((7)(−4)−5) =(2)((7)(−4)+−5) =(2)((7)(−4))+(2)(−5) =−56−10 =−66

OpenStudy (ineedhelplz):

Thats what I got...

OpenStudy (ayonnaleflore):

when i did the problem i got -231

OpenStudy (ineedhelplz):

hmm where is your work?

OpenStudy (ineedhelplz):

@ayonnaleflore

OpenStudy (ineedhelplz):

@abbycross167 What was your initial equation

OpenStudy (ayonnaleflore):

-5-2(7x4+(-5)) -7(28=(-5)) -7(33) -231

OpenStudy (ineedhelplz):

That you worked of off

OpenStudy (ineedhelplz):

@ayonnaleflore That is incorrect...

OpenStudy (abbycross167):

b – a(cd + b) Is this what you wanted?

OpenStudy (ineedhelplz):

OH! I didn't see the B

OpenStudy (ineedhelplz):

Let me re-work my problem

OpenStudy (ineedhelplz):

and I was asking with the substituted numbers! :)

OpenStudy (ayonnaleflore):

-5-2(7x4+9(-5)) -7(28+(-5)) -7(33) -231 im still getting the same answer

OpenStudy (ineedhelplz):

−5−2((7)(−4)−5) =61

OpenStudy (ineedhelplz):

It is because we all have different equations...

OpenStudy (ayonnaleflore):

i know why im wrong......its -4 -5-2((7)(-4)-5 =61

OpenStudy (abbycross167):

Ok so The answer is just −5−2((7)(−4)−5) =61 ?

OpenStudy (ayonnaleflore):

hold on i just got a positive .......

OpenStudy (ayonnaleflore):

231......

OpenStudy (abbycross167):

@Ineedhelplz could help me some more?

OpenStudy (abbycross167):

@ayonnaleflore I think you might be doing it wrong... Because my teacher didn't say that 61 was wrong, so i think the answer is 61

OpenStudy (ayonnaleflore):

@abbycross167 i just asked my teacher and she said it was right so idk what your teacher is telling you

OpenStudy (anonymous):

-66 is correct

OpenStudy (abbycross167):

oh... hmmm this is confusing

OpenStudy (abbycross167):

thank you everyone for helping me!! :)